OpenStudy (anonymous):

A man had 3 sons and 17 Sheeps. When the man died, said before dying that his sheeps will be divided between 3 sons as 1/2 ,1/3 and 1/9 respectively to elder ,middle and youngest son. How will they didvide the sheeps without killing any sheep ?

OpenStudy (anonymous):

no of sheeps are 17. add 1 to it then it becomes 18. now according to the problem \[18*1/2=9,18*1/3=6,18*1/9=2\]so the son will get 9,3,2,sheeps respectively. yoy can cheek 9+6+2=17.

OpenStudy (anonymous):

correct answer.